On an average 5% of the population in an area suffers from Diabetes. What is the probability that out of 7 persons chosen at random from this area with less than and equals to three suffer from the disease


1
Expert's answer
2022-05-09T16:19:14-0400

Let "X=" the number of persons suffer from the disease :"X\sim Bin(n, p)."

Given "n=7 p=0.05, q=1-p=0.95."


"P(X\le)=P(X=0)+P(X=1)"

"+P(X=2)+P(X=3)"

"=\dbinom{7}{0}(0.05)^{0}(0.95)^{7-0}+\dbinom{7}{1}(0.05)^{1}(0.95)^{7-1}"

"\approx0.0002927"

"+\dbinom{7}{2}(0.05)^{2}(0.95)^{7-2}+\dbinom{7}{3}(0.05)^{3}(0.95)^{7-3}"

"=0.99980642188"